There are 13 more nickels than dimes and it equals 3.20
Answer by stanbon(75887)

There are 13 more nickels than dimes and it equals 3.20
-----------
Let number of dimes be "x"; value of these is 10x cents.
Number of nickels is "x+13" ; value of these is 5(x+13)=5x+65 cents
-----------
EQUATION:
value + value = 320 cents
10x + 5x+65= 320
15x= 255
x= 17 (number of dimes)
x+13 = 30 (number of nickels)
Cheers,
Stan H.